# Emergence of a Fermi-surface in the current-driven Hidden state of 1T-TaS2

### v1

@@ III Summary
Our study reveals that the hidden state in 1T-TaS2 exhibits a hole-like band crossing the Fermi level, resulting in a Fermi surface reminiscent of the NC-CDW phase. This similarity is not unexpected, given the structural resemblance between the two phases, both characterized by domains of commensurate CDW. However, the underlying mechanisms rendering these phases metallic remain elusive. While it was hypothesized that conducting domain walls, where the order parameter , contribute to the conductivity by forming channels, our ARPES measurements of the hidden phase suggest a different scenario.
We demonstrate that the hidden Fermi surface displays distinct chirality, originating from the C-CDW arrangements—a property inherent to the domains rather than the domain walls. Additionally, ARPES, which is sensitive to the surface area, indicates that the contribution from domain walls is negligible. This observation implies that each domain is intrinsically metallic and possesses an in-plane C-CDW configuration, suggesting that additional factors influence the metallic behavior.
We propose that the metallic behavior originates from the out-of-plane electronic dispersion. The periodicity of the dispersion along k⟂ indicates a doubling of the Brillouin zone in the out-of-plane direction within the C-CDW state, consistent with the previously reported layer dimerization Stahl et al. (2020). Upon transitioning to the hidden state, the out-of-plane dispersion changes, suggesting that star-of-David clusters on neighboring planes are no longer correlated. This leads to a profound change in the electronic structure: instead of having two electrons per unit cell in the C-CDW state, we now have a metallic half-filled band. The resulting band structure is unusual for a TMD, as it has a larger hopping constant normal to the Ta planes than within the planes.
IV Methods
IV.1 Sample preparation
@@
1T-TaS2 single crystal were prepared as described in Ref. Ribak et al. (2017). Samples of 1T-TaS2 were prepared for nanoARPES experiments. Single crystals were cut into narrow bridges 200-500 in width. The samples were glued to a sapphire substrate and 4 contacts were made using Ag epoxy. The samples were cooled to 40K and cleaved under UHV conditions. The 4 contacts were used to drive a DC current pulse and to monitor the sample resistance. We emphasis that during the ARPES measurement no current was applied to the sample.
IV.2 ARPES
@@
Spatially resolved ARPES experiments were carried out in the I05 beamline at Diamond light source, UK. The spatial resolution is set by the beam spot-size, which in our experiment was either 500nm using a Zone-plate or 5 micrometers using a capillary mirror. Energy resolution was set to about 50meV. Data in the NC-CDW state was measured at the CASSIOPEE beamline at Soleil, France. Measurements were done at room temperature using 80eV photon energy. Beam spot size was 100 microns. Energy resolution was set to 20meV. The Fermi-level was determined by measuring a metallic sample at the same conditions.
IV.3 calculation of k⟂
@@
We have used the three-step-model for the out-of-plane momentum k Damascelli (2004), where is the inner potential. We have used , obtained by following the normal-crystal periodicity over 3 BZs, demanding ().
V Acknowledgments
@@
We thank Assa Auerbach and Daniel Podolsky for useful discussions. We acknowledge Diamond Light Source for time on Beamline i05 under Proposal SI33131. We acknowledge SOLEIL for provision of synchrotron radiation facilities under proposal 20230771 and we would like to thank Chiara Bigi and Francois Bertran for assistance in using beamline CASSIOPEE. The work at the Technion was supported by Israeli Science Foundation grant number ISF-1263/21.

### v2

@@ Discussion
Our study reveals that the hidden state in 1T-TaS2 hosts a hole-like band crossing the Fermi level, resulting in a Fermi surface reminiscent of the nearly commensurate NC-CDW phase. This similarity is not surprising, given the structural resemblance between the two phases—both characterized by domains of commensurate CDW order. However, the mechanisms underlying the metallicity of these phases remain elusive. While it has been proposed that conducting domain walls—regions where the order parameter - form channels that contribute to the overall conductivity, our ARPES measurements suggest that this mechanism alone cannot account for the observed metallic behavior. Although such channels may be locally conductive, they are insufficient to explain the emergence of a robust Fermi surface and substantial spectral weight at the Fermi level.
We show that the hidden Fermi surface exhibits distinct chirality arising from the C-CDW domain arrangements - an intrinsic property of the domains themselves rather than the domain walls. Furthermore, the significant spectral weight observed in ARPES, which is surface-sensitive, indicates that each domain is intrinsically metallic, and the contribution from domain walls is minimal. Importantly, the in-plane CDW motif remains essentially intact across both insulating and metallic regions, suggesting that the metallicity does not result from CDW melting or disordering, but from a more subtle structural transformation.
Our findings establish a compelling link between the stacking configuration of 1T-TaS2 and its emergent electronic properties. While the monolayer is likely a Mott insulator [37], in the bulk, first-principles calculations have shown that the out-of-plane stacking of star-of-David clusters plays a decisive role in shaping the in-plane electronic dispersion, modulating gap structures and bandwidths depending on the stacking configuration [34, 45]. Specifically, DFT calculations by Lee et al. [34] indicate that the AL-stacking configuration is energetically the most favorable, with the L-stacking only slightly higher in energy. This small energy difference suggests that a stacking rearrangement from AL to L can be induced by a moderate external perturbation such as a current pulse.
To uncover the nature of this transformation, we examined the out-of-plane band dispersion at binding energies where differences between metallic and insulating regions are minimized, isolating the influence of interlayer effects. In the insulating regions, the kz dispersion shows a doubled periodicity and reduced bandwidth, consistent with interlayer dimerization and a doubling of the unit cell along the -axis, as previously reported [35]. These features align with the AL-stacking configuration, in which vertical dimers of star-of-David clusters are stacked such that the central atom of one dimer aligns with one of the outer atoms of a neighboring cluster (See Fig. 1B). The dashed white line in Fig.5B shows the calculated out-of-plane dispersion for this configuration, which matches well with our data from the insulating regions.
In contrast, the metallic regions exhibit a single-period kz dispersion closely matching the primitive lattice periodicity and a broader bandwidth, indicating the absence of dimerization, in agreement with the DFT predictions for the L-configuration (See Fig. 1B). These observations suggest that the current pulse disrupts the AL-stacked (dimerized) configuration, breaking the interlayer correlations that stabilize the insulating state. The system then transitions into the L-stacking configuration, a nearly degenerate configuration in energy, which restores a simpler stacking periodicity and enables coherent interlayer hopping, thereby giving rise to a three-dimensional metallic state.
This hidden metallic phase represents an unusual electronic structure among transition metal dichalcogenides: the out-of-plane (interlayer) bandwidth exceeds the in-plane bandwidth. This is in agreement with transport measurements showing that the resistivity anisotropy in the NC-CDW state is close to one [46].
Such an inversion of dimensionality, where transport and electronic coherence are dominated by the out-of-plane direction, challenges conventional views of quasi-2D physics and may open new avenues for engineering anisotropic transport and manipulating quantum phases in layered materials.
Materials and Methods
Sample preparation
@@
1T-TaS2 single crystal were grown as described in Ref. [47]. Samples of 1T-TaS2 were prepared for nanoARPES experiments: Single crystals were cut into narrow bridges 200-500 in width. The samples were glued to a sapphire substrate and 4 contacts were made using Ag epoxy. The samples were cooled to and cleaved under UHV conditions. The 4 contacts were used to drive a current pulse and to monitor the sample resistance in-situ. We emphasize that during the ARPES measurement no current was applied to the sample.
ARPES measurements
@@
Spatially resolved ARPES experiments were carried out in the I05 beamline at Diamond light source, UK. The spatial resolution is set by the beam spot-size, which in our experiment was either 500 nm using a Zone plate or 5 micrometers using a capillary mirror. Energy resolution was set to about 50 m eV. Data in the NC-CDW state was measured at the CASSIOPEE beamline at Soleil, France. Measurements were done at room temperature using 80 eV photon energy. Beam spot size was 100 microns. Energy resolution was set to 20 m eV. The Fermi-level was determined by measuring a metallic sample at the same conditions.
calculation of k⟂
@@
We have used the three-step-model for the out-of-plane momentum k [48], where is the inner potential. We have used , obtained by following the normal-crystal periodicity over 3 BZs, demanding ().
Band fitting
Peak fitting was done using Gaussian functions. When close to , the Gauassians were multiplied by the Fermi function, where was fixed to zero, and the temperature was set to .
Acknowledgments
We thank Assa Auerbach and Daniel Podolsky for useful discussions.
We acknowledge Diamond Light Source for time on Beamline i05 under Proposal SI33131.
@@ Acknowledgments
We acknowledge SOLEIL for provision of synchrotron radiation facilities under proposal 20230771 and we would like to thank Chiara Bigi and Francois Bertran for assistance in using beamline CASSIOPEE.
Funding:
The work at the Technion was supported by Israeli Science Foundation grant number ISF-1263/21.
Author contributions:
A.K. conceived the experiment. I.F. prepared the 1T-TaS2 single crystals. Y.N. performed the ARPES experiments and analyzed the ARPES data. R.A.G.K, I.M., A.D., N.R. S.K.P. and Y.J. performed ARPES measurements. A.L., M.D.W. and C.C. provided support at the synchrotron beam-line. Y.N. and A.K. wrote the manuscript with contributions from all authors.
Competing interests:
The authors declare that they have no competing interests.
Data availability:
All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
Supplementary materials
Figures S1 to S3